# Text-Queried Audio Source Separation via Hierarchical Modeling

### v1

@@ Abstract
Target audio source separation with natural language queries presents a promising paradigm for extracting arbitrary audio events through arbitrary text descriptions. Existing methods mainly face two challenges, the difficulty in jointly modeling acoustic-textual alignment and semantic-aware separation within a blindly-learned single-stage architecture, and the reliance on large-scale accurately-labeled training data to compensate for inefficient cross-modal learning and separation. To address these challenges, we propose a hierarchical decomposition framework, HSM-TSS, that decouples the task into global-local semantic-guided feature separation and structure-preserving acoustic reconstruction. Our approach introduces a dual-stage mechanism for semantic separation, operating on distinct global and local semantic feature spaces. We first perform global-semantic separation through a global semantic feature space aligned with text queries. A Q-Audio architecture is employed to align audio and text modalities, serving as pretrained global-semantic encoders. Conditioned on the predicted global feature, we then perform the second-stage local-semantic separation on AudioMAE features that preserve time-frequency structures, followed by acoustic reconstruction. We also propose an instruction processing pipeline to parse arbitrary text queries into structured operations, extraction or removal, coupled with audio descriptions, enabling flexible sound manipulation. Our method achieves state-of-the-art separation performance with data-efficient training while maintaining superior semantic consistency with queries in complex auditory scenes.
Index Terms:
text-queried audio source separation, hierarchical modeling, audio representation learning.
@@ I Introduction
Real-world environmental sounds typically comprise diverse audio events from multiple sources. Target sound separation, which isolates specific sound components from mixtures across domains like speech [1, 2, 3], general audio [4], and music [5], conventionally relies on single-source training samples and focuses on separating predefined source types [6]. Recent advances in universal sound separation (USS) [7] have expanded this capability to arbitrary sound sources in real-world recordings. However, the inherent complexity of separating overlapping audio events with varying characteristics persists as a fundamental challenge. Text-queried target sound extraction (TSE), which uses natural language descriptions to selectively separate sounds, has emerged as a promising solution. Unlike audio-visual [8, 9] and audio-queried [7, 10, 11] methods, it provides greater flexibility in query formulation. It also outperforms label-queried [12, 13, 14] methods by eliminating the need for fixed label categories, thereby supporting queries of any type and facilitating open-domain generalization.
@@
Recently, an increasing number of studies have explored this task, which can mainly be divided into two categories: mask-based separators [15, 16, 17, 4] and conditional generative models [18, 19, 20]. The dominant mask-based approaches typically employ separation networks to estimate a multiplicative time-frequency mask through conditioned U-Net architectures. These methods incorporate a conditioning signal into the audio U-Net with a query network to predict the target mask. The other stream, the emerging generative paradigm by transformer [21] or diffusion [22] models, formulates separation as a specialized audio editing task [18].
@@
Despite their varied frameworks, these methods all employ a joint optimization of cross-modal understanding and acoustic separation within a single-stage end-to-end framework (shown in Figure 1 (a)). This predominantly data-driven approach could potentially lead to training instability and an increased risk of overfitting, necessitating the use of large-scale audio-text datasets. The fundamental limitations persist when operating on the spectrograms [4] or VAEs’ latent [18] representations, as they lack explicit semantics. This forces models to resolve ambiguities through acoustic details alone, often resulting in under-separation or inconsistent audio events. Furthermore, commonly used weakly-labeled datasets like AudioSet [23] and VGGSound [24] contain significant label noise and irrelevant audio content, which can confuse the trained models and further complicate text instruction understanding.
@@
In this work, we decompose the text-queried audio separation into semantic-feature-level manipulation at different levels followed by acoustic reconstruction. Different from acoustic representations (e.g., spectrograms or VAE latents), we attempt to operate separation in a more compact global-local semantic feature space. The global semantic space aligns audio and text, inside which we can extract target semantic hints. In the local semantic space that preserves not only event semantics but also local time-frequency details for reconstruction, we can extract more fine-grained semantic features conditioned on previous-stage predicted global hints. Such a fine-grained feature is then utilized to reconstruct the waveform through a generative model. This hierarchical modeling with dual-stage separation (shown in Figure 1 (b)) effectively separates target audio events that align well with the text query without using large weakly-labeled datasets.
@@
    We propose a hierarchical modeling paradigm, HSM-TSS, to separate audios on different feature levels queried by text. The decomposition of global and fine-grained local semantic representations relieves the burden of learning audio-text alignment and extracting semantically-correct audio events simultaneously, effectively increasing the data efficiency and our model generalizability.
@@
    We pretrain a text-audio aligned audio representation, Q-Audio, through contrastive learning, which outperforms the commonly used CLAP [25, 26] in several benchmarks.
@@
    We design an instruction parsing pipeline with large language models (LLMs) to turn arbitrary textual queries into structured separation types, extraction or removal, and targeted audio descriptions, better facilitating query understanding with bidirectional sound manipulation.
  4. 4.
    We conduct experiments on several source separation benchmarks and demonstrate that the proposed hierarchical modeling achieves state-of-the-art performance by both signal-level and semantic-level metrics.
@@ II-A Unconditional sound separation
Unconditional audio separation aims to decompose mixed audio signals into individual sources without relying on external queries or prior knowledge. Early research primarily focuses on domain-specific tasks, such as speech separation [6, 3] and music source separation [27]. These methods address label ambiguity problem by permutation invariant training (PIT) [28] that permutes predictions to match target signals during training, and often require post-processing to select target sources from separated outputs. However, they typically rely on single-source training data, limiting the scalability to real-world mixtures. To mitigate this, mixture invariant training (MixIT) [29] introduces unsupervised learning with multi-source mixtures. However, it tends to over-separate and requires post-selection. Subsequent work combines MixIT with pre-trained sound classifiers [30], yet these classifiers still require single-source annotations. Other approaches, like MixPIT [31], attempted direct prediction from mixtures but encountered under-separation problems. Weakly supervised methods [32] utilized sound classifiers but were constrained by fixed label categories. These limitations underscore the challenge of achieving open-domain separation without single-source supervision or post-processing.
@@ II-B Text-queried sound separation
Text-queried audio separation leverages natural language descriptions or labels to guide the extraction of target sounds from mixtures. Early methods explored label-queried separation [12, 13], which relies on predefined class labels and struggles with generalization to unseen categories. A paradigm shift emerged with language-queried audio source separation (LASS), which allowed for flexible separation using free-form text queries. Pioneering efforts like LASS-Net [15] introduced end-to-end training with audio-text pairs, while CLIPSep [16] utilized contrastive language-image pretraining (CLIP) [33] to align text and audio embeddings through a visual modality, achieving zero-shot separation. Recent advancements have incorporated multi-modal supervision [34] and hybrid training [35] to tackle data scarcity. For instance, AudioSep [4] scaled training with large audio-language datasets and showcased robust open-domain generalization. Despite these advancements, challenges remain in handling linguistic diversity (e.g., paraphrased descriptions) and noisy real-world data. Alternative approaches, like audio-visual separation [36, 9], used visual cues but were sensitive to occlusions or off-screen sounds. Conversely, text-based methods possess broader applicability, as natural language offers a scalable and intuitive interface for specifying target sources.
@@ II-C Audio editing with generative models
The most common approach for audio editing is to train specialized models for particular tasks, like style transfer [37, 38] or inpainting [39, 40]. Recently, some works have studied general audio editing following human instructions. AUDIT [18] designs several editing tasks involving adding, removal, inpainting, replacement, and super-resolution based on latent diffusion models. UniAudio [19] is an audio foundation model which takes adding, dropping and super-resolution as general audio editing tasks. In terms of this, target sound separation can be seen as a sub-task of instruction-based audio editing. However, they still leveraged common acoustic representations, like VAEs’ latent or neural codec tokens, and typically suffered from under-separation or over-separation problems. Instead, our approach investigates the hierarchical semantic modeling paradigm with pretrained representations at different feature levels, with which it achieves superior performance with carefully synthesized training data with accurate labels.
@@ II-D Audio representation learning
Recent advancements in self-supervised learning (SSL) have significantly shaped audio representation learning, empowering models to extract meaningful features from raw audio data without labeled supervision. These approaches can be broadly categorized into predictive, contrastive, and masked predictive modeling techniques, each addressing distinct challenges in audio processing. Predictive modeling approaches, such as autoregressive predictive coding (APC) [41, 42], have been pivotal in learning sequential audio representations by predicting future elements of a sequential audio input based on past data. In masked predictive approach, MAE-AST [43] adapted the masked autoencoders (MAEs) [44] for audios, combining discriminative and generative objectives for training. It splits audio spectrogram into patches, mask a portion of them, and use transformers to reconstruct the masked segments. The work AudioMAE [45], further investigated shifting windows and local self-attention mechanisms to enhance modeling. Unlike predictive and masked approaches, contrastive learning has largely been utilized to learn cross-modal representations such as CLIP [33] in computer vision. It has also been adapted for audio , as seen in CLAP [26, 25], which unifies general audio and text into a joint embedding space. These multi-modal approaches provide weak supervision for tasks like multi-source sound separation, further enriching audio representation learning. Motivated by BLIP-2 [46] that leverages a Q-Former for vision-language representation learning, we introduce a lightweight module Q-Audio as a bridge between text and audio with frozen audio and text encoders in our hierarchical modeling.
@@ III The proposed framework
Figure 2: Overview of our proposed hierarchical modeling and separation frameworks. (a) Two-level audio feature representation. (b) Text-queried two-level separation.
@@ III-A Overview
$L$ is the audio length. The goal is to separate the target audio source $x_{tgt}$ based on a given text instruction $\mathcal{T}$. This separation process can be formalized as $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto x_{tgt}$, where $\mathcal{H}$ denotes the separation model. Conventional TSE methods [15, 16, 17, 4] typically rely on a purely data-driven approach through a single-stage end-to-end blind learning architecture, where semantic and acoustic information are entangled in the mixture and it struggles to distinguish target audio events from the interference. Inspired by audio and speech generative models [22, 47], we reformulate the TSE task as a regeneration process by $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto S\mapsto x_{tgt}$. Here, an intermediate feature $S$ is introduced to decouple semantic separation with acoustic reconstruction. Such an intermediate representation $S$ is both semantic and information-rich in containing local semantics to reconstruct the target audio. We take it as local semantic feature in our approach.
To provide more robust semantic guidance, we further decompose the semantic feature $S$ by extracting a semantic-only global feature $G$ on top of it, which captures high-level audio event descriptions (e.g., dog barking) without spatial or acoustic details, and aligns well with the text feature space. Comparatively, $S$ is more fine-grained, which integrates both semantic representation (e.g. what happened in the audio) and acoustic properties (e.g., pitch or spatial positioning) necessary for waveform reconstruction. This motivates a hierarchical framework that divides the feature-level separation into global-semantic and local-semantic stages, with the predicted global feature $\hat{G}$ acting as a conditioning input to refine the extraction of the local feature $\hat{S}$. The decomposition process can be denoted as $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto G\mapsto S\mapsto x_{tgt}$.
@@
$\mathcal{H}_{1}$, $\mathcal{H}_{2}$, $\mathcal{H}_{3}$ denote global-semantic separation, local-semantic separation and acoustic decoder, respectively. As depicted in Figure 2 (b), each stage yields a specialized representation tailored to its role in the separation pipeline. The acoustic decoder maps the local-semantic feature $\hat{S}$ to the acoustic domain and then reconstructs the target waveform using a neural codec decoder. In the following sections III-C, III-D and III-E, we will describe these three modules in detail.
To enhance the model’s ability to interpret diverse instructions, we leverage a pre-trained LLM to decompose the text input into two components: the task type $T_{task}$ and the caption $T_{cap}$, as shown in the input of Figure 2. By isolating the captions from task types, we can exploit the strengths of audio-language contrastive learning that aligns text and audio in a shared representation space. We define two separation tasks: removal and extraction, allowing the captions to describe either the audio events to be removed or those to be retained, thereby increasing the flexibility of user instructions. This process is elaborated in Section III-F.
Compared to single-stage methods, this hierarchical framework offers several advantages. It disentangles semantic and acoustic processing to mitigate the impact of interference, enables modular optimization of each stage, and improves robustness in scenarios where target and interfering audio events overlap significantly. By integrating instruction decomposition with staged feature extraction, our approach achieves precise and flexible text-queried sound separation.
III-B Hierarchical audio representations

### v2

@@ Abstract
Target audio source separation with natural language queries presents a promising paradigm for extracting arbitrary audio events through arbitrary text descriptions. Existing methods mainly face two challenges, the difficulty in jointly modeling acoustic-textual alignment and semantic-aware separation within a blindly-learned single-stage architecture, and the reliance on large-scale accurately-labeled training data to compensate for inefficient cross-modal learning and separation. To address these challenges, we propose a hierarchical decomposition framework, HSM-TSS, that decouples the task into global-local semantic-guided feature separation and structure-preserving acoustic reconstruction. Our approach introduces a dual-stage mechanism for semantic separation, operating on distinct global and local semantic feature spaces. We first perform global-semantic separation through a global semantic feature space aligned with text queries. A Q-Audio architecture is employed to align audio and text modalities, serving as pretrained global-semantic encoders. Conditioned on the predicted global feature, we then perform the second-stage local-semantic separation on AudioMAE features that preserve time-frequency structures, followed by semantic-to-acoustic reconstruction. We also split text queries into structured operations, extraction or removal, coupled with audio descriptions, enabling bidirectional sound manipulation. Our method achieves state-of-the-art separation performance with data-efficient training while maintaining superior semantic consistency with queries in complex auditory scenes.
@@ I Introduction
Real-world environmental sounds typically comprise diverse audio events from multiple sources. Target sound separation, which isolates specific sound components from mixtures across domains like speech [divide_conquer, casa, luo2019conv], general audio [liu2024separate], and music [Kong2021DecouplingMA], conventionally relies on single-source training samples and focuses on separating predefined source types [wang2018supervised]. Recent advances in universal sound separation (USS) [kong2023universal] have expanded this capability to arbitrary sound sources in real-world recordings. However, the inherent complexity of separating overlapping audio events with varying characteristics persists as a fundamental challenge. Text-queried target sound extraction (TSE), which uses natural language descriptions to selectively separate sounds, has emerged as a promising solution. Unlike audio-visual [zhao2018sound, tzinis2020into] and audio-queried [kong2023universal, chen2022zero, gfeller2021one] methods, it provides greater flexibility in query formulation. It also outperforms label-queried [veluri2023real, delcroix2022soundbeam, wang2022improving] methods by eliminating the need for fixed label categories, thereby supporting queries of any type and facilitating open-domain generalization.
@@
Recently, an increasing number of studies have explored this task, which can mainly be divided into two categories: mask-based separators [liu2022separate, dong2022clipsep, mahmud2024weakly, liu2024separate, cheng2024omnisep] and conditional generative models [wang2023audit, yang2023uniaudio, han2024instructme, yuan2025flowsep]. The dominant mask-based approaches typically employ separation networks to estimate a multiplicative time-frequency mask through conditioned U-Net architectures. These methods incorporate a conditioning signal into the audio U-Net with a query network to predict the target mask. The other stream, the emerging generative paradigm by transformer [kreuk2022audiogen], diffusion [wang2023audit] or flow matching [yuan2025flowsep] models, formulates separation as a specialized audio editing task [wang2023audit].
@@
Despite their varied frameworks, these methods all employ a joint optimization of cross-modal understanding and acoustic separation within a single-stage end-to-end framework (shown in Figure 1 (a)). This predominantly data-driven approach could potentially lead to an increased risk of overfitting, necessitating the use of large-scale audio-text datasets. The fundamental limitations persist when operating on spectrograms capturing acoustic information [liu2024separate] or VAEs’ latent [wang2023audit] feature, as they lack explicit semantic representation [liu2024audioldm]. This forces models to resolve ambiguities through acoustic details alone, often resulting in under-separation or inconsistent audio events. Furthermore, commonly used weakly-labeled datasets like AudioSet [gemmeke2017audio] and VGGSound [chen2020vggsound] contain significant label noise and irrelevant audio content, which can confuse the trained models and further complicate text instruction understanding.
@@
In this work, we decompose the text-queried audio separation into semantic-feature-level manipulation at different levels followed by semantic-to-acoustic decoding and acoustic reconstruction. Different from acoustic representations (e.g., spectrograms or VAE latents), we attempt to operate separation in a more compact global-local semantic feature space. The global semantic space aligns audio and text, inside which we can extract target semantic hints. In the local semantic space that preserves not only event semantics but also local time-frequency details for reconstruction, we can extract more fine-grained semantic features conditioned on previous-stage predicted global hints. Such a fine-grained feature is then utilized to reconstruct the waveform through a semantic-to-acoustic generative model. This hierarchical modeling with dual-semantic-stage separation and semantic-to-acoustic reconstruction (shown in Figure 1 (b)) effectively separates target audio events that align well with the text query without using large weakly-labeled datasets. In addition, we formulate target audio source separation as two complementary tasks, enabling the extraction of target sources and the removal of interferences. For instance, the query “Remove the dog barking” can achieve the same goal as “Extract the rain sound” in a mixture if “rain” is the desired sound.
In this paper, semantic refers to high-level interpretations of audio and text, such as sound event categories or descriptive labels. global semantic features capture holistic, event-level information like the overall context or type of sound, without acoustic details such as volume or timbre. In contrast, local semantic features represent more fine-grained, time-frequency-aware detailed semantic information, such as specific words in speech or the timing and location of sound events, capable to reconstruct the waveform.
@@
    We propose a hierarchical modeling paradigm, HSM-TSS, to separate audios on different feature levels queried by text. By decomposing global and local semantic representations with decoupled semantic separation and acoustic reconstruction, the proposed method alleviates the challenge of learning audio-text alignment and extracting semantically-correct audio events simultaneously, effectively enhancing our model generalizability.
@@
    We pretrain a text-audio aligned audio representation, Q-Audio, through contrastive learning, which outperforms the commonly used CLAP [wu2023large, elizalde2023clap] in several benchmarks, including AudioCaps and Clotho for both text-to-audio retrieval and audio captioning.
@@
    We conduct experiments on several datasets including Clotho, AudioCaps, and FSD50K and demonstrate that the proposed hierarchical modeling achieves state-of-the-art separation performance by both signal-level and semantic-level metrics.
@@ II-A Unconditional sound separation
Unconditional audio separation aims to decompose mixed audio signals into individual sources without relying on external queries or prior knowledge. Early research primarily focuses on domain-specific tasks, such as speech separation [wang2018supervised, luo2019conv] and music source separation [takahashi2021densely]. These methods address label ambiguity problem by permutation invariant training (PIT) [yu2017permutation] that permutes predictions to match target signals during training, and often require post-processing to select target sources from separated outputs. However, they typically rely on single-source training data, limiting the scalability to real-world mixtures. To mitigate this, mixture invariant training (MixIT) [wisdom2020unsupervised] introduces unsupervised learning with multi-source mixtures. However, it tends to over-separate and requires post-selection. Subsequent work combines MixIT with pre-trained sound classifiers [wisdom2021sparse], yet these classifiers still require single-source annotations. Other approaches, like MixPIT [karamatli2022mixcycle], attempted direct prediction from mixtures but encountered under-separation problems. Weakly supervised methods [pishdadian2020finding] utilized sound classifiers but were constrained by fixed label categories. These limitations underscore the challenge of achieving open-domain separation without single-source supervision or post-processing.
@@ II-B Text-queried sound separation
Text-queried audio separation leverages natural language descriptions or labels to guide the extraction of target sounds from mixtures. Early methods explored label-queried separation [veluri2023real, delcroix2022soundbeam], which relies on predefined class labels and struggles with generalization to unseen categories. A paradigm shift emerged with language-queried audio source separation (LASS), which allowed for flexible separation using free-form text queries. Pioneering efforts like LASS-Net [liu2022separate] introduced end-to-end training with audio-text pairs, while CLIPSep [dong2022clipsep] utilized contrastive language-image pretraining (CLIP) [radford2021learning] to align text and audio embeddings through a visual modality, achieving zero-shot separation. AudioSep [liu2024separate] scaled training with large audio-language datasets and showcased robust open-domain generalization. OmniSep [cheng2024omnisep] further enabled separating target audios through both single-modal and multi-modal composed queries. Some recent advancements have leveraged audio‑only data, i.e., without paired captions, to mitigate the lack of annotations for text‑queried extraction [saijo2025leveraging, ma2025language]. Researchers have also been exploring audio separation with generative models like diffusion models [hai2024dpm] and rectified flow matching [yuan2025flowsep]. Despite these advancements, challenges remain in handling linguistic diversity (e.g., paraphrased descriptions) and noisy real-world data. Alternative approaches, like audio-visual separation [gao2019co, tzinis2020into], used visual cues but were sensitive to occlusions or off-screen sounds. Conversely, text-based methods possess broader applicability, as natural language offers a scalable and intuitive interface for specifying target sources.
@@ II-C Audio editing with generative models
The most common approach for audio editing is to train specialized models for particular tasks, like style transfer [wu2023musemorphose, popov2021diffusion] or inpainting [borsos2022speechpainter, marafioti2020gacela]. Recently, some works have studied general audio editing following human instructions. AUDIT [wang2023audit] designs several editing tasks involving adding, removal, inpainting, replacement, and super-resolution based on latent diffusion models. UniAudio [yang2023uniaudio] is an audio foundation model which takes adding, dropping and super-resolution as general audio editing tasks. In terms of this, target sound separation can be seen as a sub-task of instruction-based audio editing. However, they still leveraged common acoustic representations, like VAEs’ latent or neural codec tokens, and typically suffered from under-separation or over-separation problems. Instead, our approach investigates the hierarchical semantic modeling paradigm with pretrained representations at different feature levels, with which it achieves superior performance with carefully synthesized training data with accurate labels.
@@ II-D Audio representation learning
Recent advancements in self-supervised learning (SSL) have significantly shaped audio representation learning, empowering models to extract meaningful features from raw audio data without labeled supervision. These approaches can be broadly categorized into predictive, contrastive, and masked predictive modeling techniques, each addressing distinct challenges in audio processing. Predictive modeling approaches, such as autoregressive predictive coding (APC) [chung2020generative, chung2020vector], have been pivotal in learning sequential audio representations by predicting future elements of a sequential audio input based on past data. In masked predictive approach, MAE-AST [baade2022mae] adapted the masked autoencoders (MAEs) [he2022masked] for audios, combining discriminative and generative objectives for training. It splits audio spectrogram into patches, mask a portion of them, and use transformers to reconstruct the masked segments. The work AudioMAE [huang2022masked], further investigated shifting windows and local self-attention mechanisms to enhance modeling. Unlike predictive and masked approaches, contrastive learning has largely been utilized to learn cross-modal representations such as CLIP [radford2021learning] in computer vision. It has also been adapted for audio , as seen in CLAP [elizalde2023clap, wu2023large], which unifies general audio and text into a joint embedding space. These multi-modal approaches provide weak supervision for tasks like multi-source sound separation, further enriching audio representation learning. Motivated by BLIP-2 [li2023blip] that leverages a Q-Former for vision-language representation learning, we introduce a lightweight module Q-Audio as a bridge between text and audio with frozen audio and text encoders in our hierarchical modeling. Similar Q-Former-style bridging between speech/audio and text has also been explored in prior work, e.g. ZeroST [khurana2024zerost] for zero‑shot speech translation.
@@ III The proposed framework
Figure 2: Overview of our proposed hierarchical modeling and separation frameworks. (a) Two-level audio feature representation. (b) Text-queried two-level separation and decoupled semantic separation and acoustic reconstruction.
@@ III-A Overview
$L$ is the audio length. The goal is to separate the target audio source $x_{tgt}$ based on a given text instruction $\mathcal{T}$. This separation process can be formalized as $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto x_{tgt}$, where $\mathcal{H}$ denotes the separation model. Conventional TSE methods [liu2022separate, dong2022clipsep, mahmud2024weakly, liu2024separate] typically rely on a purely data-driven approach through a single-stage end-to-end blind learning architecture, where semantic and acoustic information are entangled in the mixture and it struggles to distinguish target audio events from the interference. Inspired by audio and speech generative models [liu2024audioldm, borsos2023audiolm], we reformulate the TSE task as a regeneration process by $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto S\mapsto x_{tgt}$. Here, an intermediate feature $S$ is introduced to decouple semantic separation with acoustic reconstruction. Such an intermediate representation $S$ is both semantic and information-rich in containing local semantics to reconstruct the target audio. We take it as local semantic feature in our approach.
To provide more robust semantic guidance, we further decompose the semantic feature $S$ by extracting a semantic-only global feature $G$ on top of it, which captures high-level audio event descriptions (e.g., dog barking) without spatial or acoustic details.Here, spatial/acoustic details refer to fine-grained characteristics of the audio signal (e.g. spatial cues or timbral nuances) that are typically not fully described in the accompanying text. This design ensures that $G$ aligns well with the text feature space. Comparatively, $S$ is more fine-grained, which integrates both semantic representation (e.g. what happened in the audio) and acoustic properties (e.g., pitch or spatial positioning) necessary for waveform reconstruction. This motivates a hierarchical framework that not only decouples semantic separation with acoustic reconstruction but also divides the feature-level separation into global-semantic and local-semantic stages, with the predicted global feature $\hat{G}$ acting as a conditioning input to refine the extraction of the local feature $\hat{S}$. The decomposition process can be denoted as $\mathcal{H}:(x_{mix},\mathcal{T})\mapsto G\mapsto S\mapsto x_{tgt}$.
@@
$\mathcal{H}_{1}$, $\mathcal{H}_{2}$, $\mathcal{H}_{3}$ denote global-semantic separation, local-semantic separation and semantic-to-acoustic and acoustic decoders, respectively. As depicted in Figure 2 (b), each stage yields a specialized representation tailored to its role in the pipeline. The semantic-to-acoustic and acoustic decoder maps the local-semantic feature $\hat{S}$ to the acoustic domain and then reconstructs the target waveform using a neural codec decoder. In the following sections III-C, III-D and III-E, we will describe these three modules in detail.
We decompose the text input into two components: the task type $T_{task}$ and the caption $T_{cap}$, as shown in the input of Figure 2. By isolating the captions from task types, we can exploit the strengths of audio-language contrastive learning that aligns text and audio in a shared representation space. We define two separation tasks: removal and extraction, allowing the captions to describe either the audio events to be removed or those to be retained. This process is elaborated in Section III-F.
Compared to single-stage methods, this three-level hierarchical framework offers several advantages. It disentangles semantic and acoustic processing to mitigate the impact of interference, enables modular optimization of each stage, and improves robustness in challenging scenarios where target and interfering audio events overlap significantly as evidenced in the experimental part. By integrating instruction decomposition with staged feature extraction, our approach achieves precise and flexible text-queried sound separation.
III-B Two-level semantic representations